# Vamos: Versatile Action Models for Video Understanding

### v1

@@ 4.9 Comparison with State-of-the-art
Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 678 and 9. On EgoSchema zero-shot VQA, our approach based on GPT-4 outperforms the best vision-language model by 50.3%. On NeXT-QA, Vamos achieves the best performance among baselines with similiar model size. It is slightly worse than SeViLA, which trains on additional dataset for temporal localization. On IntentQA, Vamos outperforms all baselines, with a 17.4% accuracy improvement. On Ego4d LTA, Vamos outperforms AntGPT without relying on domain-specific video encoders [47].
Model Input Type Frames
VIOLET [18] frame 19.9%
FrozenBiLM [75] frame 26.9%
mPLUG-Owl [77] frame 31.1%
InternVideo [68] frame 32.1%
Llama2-13B caption 36.73%
GPT-3.5 caption 41.24%
GPT-4 caption 48.26%
Table 6: Egoschema VQA zero-shot Performance.
@@
Vamos (ours) 72.6% 69.6% 78.0% 72.5%
SeViLA[79]* 74.2% 69.4% 81.3% 73.8%
Table 7: Comparison on NeXT-QA benchmark. * denotes model using additional dataset during training.
@@
Vamos (ours) 69.52% 70.19% 64.95% 68.51%
Table 8: Comparison with SOTA on IntentQA
@@
Table 9: Comparison with SOTA on Ego4D v2 test set
@@ 5 Conclusion
We study the form of representation for long videos and proposed versatile action models (Vamos) for video understanding. Through extensive experiments on long-term action anticipation and video question answer benchmarks, we have demonstrated that surprisingly, direct application of free-form text descriptions generated by vision-language models serve as strong video representation for all benchmarks we consider, and the use of large language models enable us to perform zero-shot reasoning, and correct the answers via causal intervention. Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, close to SOTA performance on NeXT-QA, and outperforms the best vision-language model by over 50% on EgoSchema. Although our results show the promise of free-form text representation, we believe visual information is still essential for complex video understanding and reasoning, and that further explorations on modeling and benchmarking are needed.
Acknowledgements: This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology. We would like to thank Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their generous feedback on this work. Our research was conducted using computational resources at Brown CCV.

### v2

@@ 4.7 Comparison with State-of-the-art
Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 4.74.74.7, and 4.7. We train Vamos with LLaMA2-7B on Ego4D LTA and LLaMA3-8B on NeXT-QA and IntentQA. On EgoSchema zero-shot VQA, our approach based on GPT-4o outperforms the best vision-language model by 66.8% on the full set. On NeXT-QA, Vamos achieves the best performance and significantly outperforms LLaMA-VQA [36] and SeViLA[83], even though the former uses a LLM with 33B parameters and the latter is trained on additional dataset with temporal localization supervision. On IntentQA, Vamos also outperforms all baselines, with a 28.7% accuracy improvement compared to the best performing prior method. On Ego4D LTA, Vamos outperforms previous works, without relying on domain-specific video encoders [47].
Table 7: Comparison on NeXT-QA benchmark. * with additional supervision.
@@
LLaMA-VQA-33B [36] 76.2% 72.6% 78.8% 75.5%
Vamos (ours) 77.2% 75.3% 81.7% 77.3%
Table 8: Comparison with SOTA on IntentQA
@@
Vamos (ours) 75.14% 77.44% 69.58% 74.16%
Table 9: Comparison with SOTA on Ego4D LTA v2 test set.
@@
Table 10: Egoschema VQA zero-shot performance on full set and subset.
Model Input Type Full Subset
FrozenBiLM [79] frame 26.9% -
mPLUG-Owl [81] frame 31.1% -
InternVideo [71] frame 32.1% -
Vamos (LLaMA2-13B) caption 36.73% 38.20%
Vamos (GPT-3.5) caption 41.24% 47.60%
Vamos (GPT-4) caption 48.26% 51.20%
Vamos (GPT-4o) caption 53.55% 57.20%
@@ 5 Conclusion
We study different forms of video representations and propose versatile action models (Vamos) as a unified framework to utilize visual- and text-based representations for video understanding. We conduct extensive experiments on long-term action anticipation and video question answer benchmarks. Surprisingly, we observe that direct applications of free-form, general-purpose text-based video representations, such as captions, serve as strong video representation for all benchmarks we consider. Vamos utilizes large language models to perform zero-shot reasoning, and incorporate human feedback via test-time intervention. We further propose the token bottleneck models, which allow the users to interpret the evidence selected by Vamos, and speed up inference by nearly 5x. Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, NeXT-QA, and outperforms the best vision-language model by over 66% on EgoSchema.
Limitations: Although our results show the promise of free-form text-based representations, we believe visual information is still essential for complex video understanding and reasoning. We expect future work to investigate alternative visual encoders [66] to extract fine-grained visual information beyond what has been captured by the captions, and to propose pre-training paradigms that better align visual inputs with the input space of LLMs. We also believe better benchmarks that require fine-grained, structured visual understanding are needed to rigorously evaluate the impact of representations for video understanding.
Acknowledgements: This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology. We would like to thank Karttikeya Mangalam, Raiymbek Akshulakov, and Shyamal Buch for their kind help with EgoSchema and ATP; Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their useful feedback and discussions. Our research was conducted using computational resources at the Center for Computation and Visualization at Brown University.